OpenStudy (anonymous):

What is the 25th term of the arithmetic sequence where a1 = 8 and a9 = 48 ?

OpenStudy (amistre64):

a1 = 8 a2 = a1 + d a3 = (a1 + d) + d a4 = ((a1+d)+d)+d = a1 + 3d an = a1 + d(n-1)

OpenStudy (amistre64):

a9 = 48 = 8 + d(7) to solve for d

OpenStudy (amistre64):

i got lost; 48 = 8 + d(8) :)

OpenStudy (amistre64):

looks like d=5 in this case a25 = a1 + d(24) a25 = 8 + 5(24)